OpenStudy (anonymous):

x+y+z=7 x-y+z=5 x-y-z=-3 solve the system. please show me how to work it out

Parth (parthkohli):

\[x + y + z = 7 \tag{1}\]\[x - y +z = 5 \tag{2}\]\[x-y-z=-3\tag{3}\]The most effective way to solve is elimination. Add \((1)\) and \((2)\).

OpenStudy (theviper):

right :)))

OpenStudy (anonymous):

I don't understand

Parth (parthkohli):

Not sure how to explain... I think you could use the online resources.

OpenStudy (saifoo.khan):

x+y+z=7 x-y+z=5 x-y-z=-3 Let's take 2 equations first. Our goal is to eliminate one variable first. x+y+z=7 x-y+z=5 Make z the subject for the first one. z = 7 - y -x

OpenStudy (saifoo.khan):

Once you have this, substitute this value of "z" in second equation. x-y+z=5 <-2 x - y + (7-y-x) = 5 Do you agree till here?

OpenStudy (anonymous):

ok yea

OpenStudy (saifoo.khan):

Simplify the new equation. x - y + (7-y-x) = 5 x - y + 7 -y -x = 5 -2y + 7 = 5 -2y = 5 - 7 y = -2/ -2 y = 1 Agree?

OpenStudy (anonymous):

ok yeah

OpenStudy (saifoo.khan):

Now again take 2 equation. Let's take the last two now. x+y+z=7 x-y-z=-3 Solve the first one for x. x = 7 -z - y Substitute this value in the second one. x - y - z = -3 (7-z-y) -y-z =-3 Afer simplification we get -2y -2z = -10

OpenStudy (saifoo.khan):

We already knew the value of y. Which was 1. -2(1) -2z = -10

OpenStudy (anonymous):

ok..

OpenStudy (saifoo.khan):

Now solve for z to get it's value.

OpenStudy (saifoo.khan):

Finally you'll have the value of y and z. Substitute it in any equation to get the value of x.

OpenStudy (anonymous):

2,1,4 ?

OpenStudy (saifoo.khan):

^

OpenStudy (anonymous):

thanks so much